OpenStudy (anonymous):

What is the value of f(4) for the function f(x) = 3x – 5?

jimthompson5910 (jim_thompson5910):

replace all copies of x with 4 and evaluate

jimthompson5910 (jim_thompson5910):

f(x) = 3x - 5 f(4) = 3*4 - 5 f(4) = ??

OpenStudy (anonymous):

f(4)=3(4)-4 = solve

OpenStudy (anonymous):

f(4)=8

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

3*4 = ???

jimthompson5910 (jim_thompson5910):

in other words, can you tell me what 3 times 4 is?

OpenStudy (anonymous):

12

OpenStudy (anonymous):

sorry was away

OpenStudy (anonymous):

so its f(4)=12?

OpenStudy (anonymous):

without variables 12

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

3 times 4 is 12 so 3*4 = 12 and that means f(4) = 3*4 - 5 turns into f(4) = 12 - 5

jimthompson5910 (jim_thompson5910):

the last thing you need to do is evaluate 12 - 5

OpenStudy (anonymous):

7

jimthompson5910 (jim_thompson5910):

good, so f(4) = 7

OpenStudy (anonymous):

with no variables the answer is 7?

jimthompson5910 (jim_thompson5910):

the final answer is 7, yep

OpenStudy (anonymous):

can you help me more?

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

sure I can help with a few more

OpenStudy (anonymous):

What is the slope of the line through the points (2, 0) and (3, -3)

jimthompson5910 (jim_thompson5910):

use the slope formula m = (y2 - y1)/(x2 - x1)

jimthompson5910 (jim_thompson5910):

so essentially you're subtracting the y coordinates from each other then the x coordinates from each other (in the same order) finally you are dividing and reducing

OpenStudy (anonymous):

slope is 3?

jimthompson5910 (jim_thompson5910):

very close

jimthompson5910 (jim_thompson5910):

double check your work though

jimthompson5910 (jim_thompson5910):

see where you went wrong?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is it -3

jimthompson5910 (jim_thompson5910):

how did you get 3? can you show me your work?

jimthompson5910 (jim_thompson5910):

it is -3, hopefully you see why this is?

OpenStudy (anonymous):

didnt use - at first thanks for confirming

jimthompson5910 (jim_thompson5910):

ah i gotcha